OpenStudy (anonymous):

Determine whether each equation has one solution, no solution, or infinite solutions show work. 3(x+5)+x=4x. 4(1-x)=5x+8. x+5+2x=5+3x

OpenStudy (kidrah69):

well first we need to solve this lets start with the first one :) \[3(x+5)+x=4x\] how would you solve this? any ideas?

OpenStudy (anonymous):

I solved the first one it 15=0 no solution

OpenStudy (kidrah69):

No soltuion is correct for that one what about #2?

OpenStudy (anonymous):

I think it has one solution

OpenStudy (kidrah69):

yep thats correct =)

OpenStudy (kidrah69):

What about the 3rd one?

OpenStudy (anonymous):

It is infinite solutions but what I need help with is how to show my work because I can't figure that out

OpenStudy (kidrah69):

ok so solving the equations =) ok number 1 can you solve that? Remeber we need to combine the like terms

OpenStudy (kidrah69):

But first there are parenthses so we must distribute

OpenStudy (anonymous):

Number one was the only one I was able to show my work on it number 2,and 3 that I need help on

OpenStudy (kidrah69):

2) 1 Expand 4−4x=5x+8 2 Subtract 4 from both sides −4x=5x+8−4 3 Simplify 5x+8−4 to 5x+4 −4x=5x+4 4 Subtract 5x from both sides −4x−5x=4 5 Simplify −4x−5x to −9x −9x=4 now u can get the answer just divide

OpenStudy (anonymous):

So I divide -9 and 4

OpenStudy (anonymous):

Is it -2.25

OpenStudy (kidrah69):

leave in fraction form :)

OpenStudy (anonymous):

So it is -9\4 or 4\-9

OpenStudy (kidrah69):

-4/9 :)

OpenStudy (kidrah69):

for 3 1 Cancel 5 on both sides x+2x=3x 2 Simplify x+2x to 3x 3x=3x 3 Since both sides equal, there are infinitely many solutions Infinitely Many Solutions

OpenStudy (anonymous):

For number two is 4\9 both negative numbers or just the 4

OpenStudy (kidrah69):

-4/9

OpenStudy (anonymous):

OK so its just the four

OpenStudy (kidrah69):

Bruh i itrally typed the answer there .-. \(\huge\color{red}{-4/9}\)

OpenStudy (kidrah69):

\(\huge\color{red}{WELCOME~TO~OPENSTUDY}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an image tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/

OpenStudy (anonymous):

Ok thank you I want to give you a medal how do I do that

OpenStudy (kidrah69):

Click the blue best repsone button and thank you!

OpenStudy (anonymous):

Thanks for all your help gnight

OpenStudy (kidrah69):

gnite